OpenStudy (anonymous):

Part 1 Is the following equation true? Part 2 Use complete sentences to explain the properties used in making your decision. 4 log 3 x + 1/2 log 3 y - 2 log 3 z = log 3 x^4 Square root of y/ z^2

OpenStudy (anonymous):

\[4 \log _{3} x +\frac{ 1 }{ 2 } \log _{3} y-2 \log _{3} z = \log _{3} \frac{ x ^{4} \sqrt{y}}{z ^{2} }\]

OpenStudy (anonymous):

@Luigi0210 think you could help me out on this one?

OpenStudy (anonymous):

itis right

OpenStudy (anonymous):

@satellite73 how would I explain it though? how did you come to that conclusion?

OpenStudy (anonymous):

for example \(4\log(x)=\log(x^4)\) because \(n\log(x)=\log(x^n)\)

OpenStudy (anonymous):

which also gives \[\frac{1}{2}\log(y)=\log(y^{\frac{1}{2}})=\log(\sqrt{y})\]

OpenStudy (anonymous):

and also \(\log(A)+\log(B)=\log(AB)\)

OpenStudy (anonymous):

so the fact that 1/2log(y)=log square root of y makes this equation true? @satellite73

OpenStudy (anonymous):

@skullpatrol any input?